OpenStudy (anonymous):

can anyone plz solve y'' - y = 8*e^-3*x ???? plz it is urgent ......

OpenStudy (shubhamsrg):

e^(-3x) or (e^-3)*x Am assuming the former one ?

OpenStudy (anonymous):

the 1st one. and also the paricular solution is e^(-3*x)..

OpenStudy (shubhamsrg):

y= C e^(-3x) works right ? y' = -3C( e^-3x) y'' = 9C(e^-3x) => 9C(e^(-3x)) - C(e^-3x) = 8(e^(-3x)) C=1 so y=e^(-3x) is a solution.

OpenStudy (anonymous):

what about the value of C ??????? can it be any cosnt or somthing else ?

OpenStudy (shubhamsrg):

We just solved it, C=1

OpenStudy (shubhamsrg):

infact, Ae^x + Ce^(-3x) can also be a soltuion.

OpenStudy (shubhamsrg):

where C=1 and A can be anything

OpenStudy (anonymous):

thnak u....

OpenStudy (shubhamsrg):

I am pretty sure there might be more solutions to it. Nevermind